OpenStudy (anonymous):

Which graph represents the inequality five times the quantity seven minus twelve x end quantity minus thirteen is less than or equal to negative eight?

OpenStudy (anonymous):

a number line shows a closed dot at two and an arrow extending from that dot pointing toward the right a number line shows a closed dot at two and an arrow extending from that dot pointing toward the left a number line shows a closed dot at one-half and an arrow extending from that dot pointing toward the right a number line shows a closed dot at one-half and an arrow extending from that dot pointing toward the left

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

\[5(7-12x)-13\le-8\] \[35-60x-13\le-8\] \[-60x+22\le-8\] \[-60x \le-30\] \[x \ge\frac{1}{2}\]

OpenStudy (anonymous):

so its the last one?

jigglypuff314 (jigglypuff314):

greater than (>) means to the right

OpenStudy (anonymous):

so its the second to last one!

jigglypuff314 (jigglypuff314):

yup